OpenStudy (anonymous):

Periodic function help? The function f(theta) and g(theta) re sine functions, where f(0) = g(0) = 0. The amplitude of f(theta) is twice the amplitude of g(theta). The period of f(theta) is one-half the period of g(theta). If g(theta) has a period of 2pi and f(pi/4) = 4, write the function rule for g(theta). Explain your reasoning. Ok, so I know the standard form is f(x) = a sin b theta (where a is amplitude and b is the period), but i'm lost on how to use this, but I know it has to have something to do with this.

OpenStudy (anonymous):

@amistre64 Can you help?

OpenStudy (anonymous):

@Hero Do you know how to solve this?

OpenStudy (perl):

if g(Θ) has a period of 2pi, then it will have the form g(Θ) = a sin(Θ)

OpenStudy (perl):

f(Θ) has period one half of g, so the period of f is π

OpenStudy (perl):

g(Θ) = a sin(Θ) f(Θ) = 2a*sin(2Θ) and you are given that f(π/4) = 4

OpenStudy (perl):

Therefore 4 = 2*a sin(π/4) can you solve it now?

OpenStudy (anonymous):

So 4 = sqrt(2)a?

OpenStudy (anonymous):

g(Θ) = a sin(2pi theta). I just need the amplitude :/

OpenStudy (perl):

right, so you can solve for `a`

OpenStudy (anonymous):

I get sqrt[8] = a? 4 = sqrt(2)a 16 = 2(a)^2 8 = a^2 sqrt[8] = a?

OpenStudy (anonymous):

@perl ?

OpenStudy (perl):

yes that's `a`

OpenStudy (perl):

which is simplified as 2√2

OpenStudy (anonymous):

Awesome! So g(theta) = 2(sqrt[2]) * (sin(2pi theta)) Thanks!

OpenStudy (perl):

no thats incorrect

OpenStudy (perl):

y = sin Θ , has a period of 2pi, you don't want to introduce another factor of 2pi in there

OpenStudy (anonymous):

Ooh, so just g(theta) = 2(sqrt[2]) * (sin theta)?

OpenStudy (perl):

g(Θ) = 2√2 sin(Θ) f(Θ) = 2(2√2) sin(2Θ)

OpenStudy (anonymous):

Awesome. Thanks! This can get a bit confusing for me, so I appreciate the patience!